If an element has multiple common oxidation states whivh one should we take to caluculate the E.A.N.
Example:Caluculate the E.A.N for [Co(NH3)6]Cl3

Grade:11

1 Answers

Gagan Deep
9 years ago
First find the oxidation state of Co in [Co(NH3)6]Cl3
Oxidation State of Clis -3 since there are three chlorine atoms
So, Oxidation State of [Co(NH3)6] is +3
Here oxidation state of NH3 is zero
So, Oxidation state of Co is +3
 
Now E.A.N of [Co(NH3)6]Clis (27-3)+(2 x 6) = 24+12 = 36